OpenStudy (anonymous):

complete the equation for the linear function whose graph contains the point (7,8) and has a slope of -1.2 I do not understand how to work this help

OpenStudy (theyankee):

Any equation has a graph (y - y1) = m(x - x1) that is a straight line. It passes through (x1, y1) and has slope m. The formula for finding the point slope form between any two points as follows Point Slope Form (y - y1) = m(x - x1) (x1, y1) are the coordinates of a known point on the line So, plug in your equation to get (y-8)=-1.2(x-7) And solve!